Bakery just paid an annual dividend of $2.40 a share and is expected to increase that amount by 2.0 percent per year. You are planning to buy 1,000 shares of this stock in two years. If investors require a 14.5% return on Friendly Bakery's stock. What is the current stock price? How much should you expect to pay in total when you buy 1000 shares in two years?
Answer #1
P0 = D0*(1+g)/(Ke-g) = 2.40*(1+2%)/(14.5%-2%) 19.584
Current stock price 19.58
P2 = D0*(1+g)^3/(Ke-g) = 2.40*(1+2%)^3/(14.5%-2%) 20.38
Stock price after two years 20.38
Amount to be paid to buy 1000 shares in two years = 1000 * 20.38 20380